Cedar Company has the following information related to a new project: Initial investment: $1,498,000; Fixed costs:...

  1. Cedar Company has the following information related to a new project: Initial investment: $1,498,000; Fixed costs: $416,000; Variable costs: $12.30 per unit; Selling price: $31.20 per unit; Discount rate: 13 percent; Project life: 5 years; Tax rate: 25 percent. Fixed assets are depreciated using straight-line depreciation over the project's life. What is the financial break-even point?

    46,773 units

    44,275 units

    48,924 units

    50,008 units

    42,683 units
Answer #1

Depreciation = Cost of Investment / Useful life years = $1,498,000 / 5 = $299,600

The financial break-even point is the number of units that must be sold to generate a NPV of 0.

First, calculate the Operating Cash Flow that results in a NPV of 0.

NPV = -$1,498,000 + [OCF * {(1 - 1.13-5) / 0.13}] = 0

OCF = [$1,498,000*0.13] / [1 - 1.13-5]

OCF = $194,740 / 0.4572 = $425,903.19

Next, determine the quantity that must be sold to achieve the calculated OCF.

Operating Cash Flow = [{Q*(Price - Variable Cost)} - Fixed Costs - Depreciation]*(1 - Tax Rate) + Depreciation

$425,903.19 = [{Q * ($31.20 - $12.30)} - $416,000 - $299,600] * (1 - 0.25) + $299,600

Q = [{($425,903.19 - $299,600)/(1-0.25)} + $416,000 + $299,600]/($31.20 - $12.30)

Q = [$168,404.25 + $715,600] / $18.90

Q = $884,004.25 / $18.90 = 46,772.71, or 46,773 units

So, 1st option is correct.